# Ask in Any Modality: A Comprehensive Survey on Multimodal Retrieval-Augmented Generation

### v2

@@ 6 Open Problems and Future Directions
Despite rapid advancements in multimodal RAG systems, fundamental challenges remain in achieving robust, efficient, and human-like reasoning across modalities.
6.1 Generalization, Explainability, and Robustness
Multimodal RAG systems often struggle with domain adaptation and exhibit modality biases, frequently over-relying on text for both retrieval and generation Winterbottom et al. (2020). Explainability remains a major challenge, as these systems typically fail to attribute answers to precise sources. Current methods often cite entire documents or large visual regions as source attribution rather than identifying the exact part of an image, speech, or other modality that led to the answer Ma et al. (2024b); Hu et al. (2023).
In addition, the interplay between modalities affects the quality of outcomes produced by these models; for example, answers derived solely from text sources may differ in quality compared to those requiring a combination of text and image inputs Baltrusaitis et al. (2019). They are also vulnerable to adversarial perturbations, such as misleading images influencing textual outputs, and their performance can degrade when relying on low-quality or outdated sources Chen et al. (2022b). While the trustworthiness of unimodal RAGs has been studied Zhou et al. (2024d), enhancing the robustness of multimodal RAGs remains an open challenge and a promising research direction.
6.2 Reasoning, Alignment, and Retrieval Enhancement
Multimodal RAGs struggle with compositional reasoning, where information from different modalities must be logically integrated to generate coherent and contextually rich outputs. While cross-modal techniques such as Multimodal-CoT Zhang et al. (2023b) have been proposed, further innovations are needed to improve the coherence and contextual relevance of multimodal outputs. Enhancing modality alignment and retrieval strategies, particularly for entity-aware retrieval, is essential. Moreover, despite the potential of knowledge graphs to enrich cross-modal reasoning, they remain largely underexplored in multimodal RAGs compared to text-based RAGs Zhang et al. (2024f); Procko and Ochoa (2024).
Retrieval biases such as position sensitivity Hu et al. (2024c), redundant retrieval Nan et al. (2024), and biases propagated from training data or retrieved content Zhai (2024), pose significant challenges that require further attention. Another promising direction is developing a unified embedding space for all modalities, enabling direct multimodal search without intermediary conversion models (e.g., ASRs). Despite some progress, mapping multimodal knowledge into a unified space remains an open challenge with significant potential.
6.3 Agent-Based and Self-Guided Systems
Recent trends indicate a shift towards agent-based multimodal RAGs that integrate retrieval, reasoning, and generation across diverse domains. Unlike static RAG systems, future multimodal RAGs should integrate interactive feedback and self-guided decision-making to iteratively refine outputs. Existing feedback mechanisms often fail to accurately determine whether errors stem from retrieval, generation, or other stages Dong et al. (2024b). The incorporation of reinforcement learning and end-to-end human-aligned feedback into multimodal RAGs remains largely unexplored but holds significant potential for enhancing these systems. These methods could enable multimodal RAGs to assess whether retrieval is necessary, evaluate the relevance of retrieved content, and dynamically determine the most suitable modalities for response generation. Achieving robust support for any-to-any modality is essential for adaptability in open-ended tasks Wu et al. (2024b).
Future multimodal RAGs should incorporate data from diverse real-world sources, such as environmental sensors, alongside traditional modalities like text and images, to enhance situational awareness. This progression aligns with the trend toward embodied AI, where models integrate knowledge with physical interaction, enabling applications in robotics, navigation, and physics-informed reasoning. Bridging retrieval-based reasoning with real-world agency brings these systems closer to AGI.
@@ 6.3 Agent-Based and Self-Guided Systems
6.4 Long-Context Processing, Efficiency, Scalability, and Personalization
High computational costs in video frame sampling and memory bottlenecks in processing multi-page documents with images remain key challenges in long-context processing. Fixed extraction rates struggle to capture relevant frames, requiring adaptive selection based on content complexity and movement Kandhare and Gisselbrecht (2024). Additionally, retrieval speed-accuracy trade-offs in edge deployments and redundant computations in cross-modal fusion layers emphasize the need for efficient, scalable architectures. Personalization mechanisms, such as adapting retrieval to user-specific contexts like medical history, remain underexplored. As personalization mechanisms evolve, ensuring privacy and mitigating the risks of sensitive data leakage in multimodal outputs remain critical challenges. Lastly, the lack of datasets with complex reasoning tasks and multimodal adversarial examples limits robust evaluation.
@@ 7 Conclusion
This study provides a comprehensive review of multimodal Retrieval-Augmented Generation (Multimodal RAG) literature. Specifically, we explore and categorize key advancements across different aspects of multimodal RAG systems, including retrieval, multimodal fusion, augmentation, generation, and training strategies. Additionally, we examine the tasks these systems address, their domain-specific applications, and the datasets, benchmarks, and evaluation methods. We also discuss open challenges and limitations in current approaches, along with promising future directions. We hope this work inspires future research, particularly in enhancing cross-modal reasoning and retrieval, developing agent-based interactive systems, and advancing unified multimodal embedding spaces.
@@ Appendix A Taxonomy
Figure 2 provides an overview of recent advances in multimodal retrieval-augmented generation (RAG) systems. The taxonomy is organized into several key categories.
  • Retrieval strategies cover efficient search and similarity retrieval methods (including maximum inner product search (MIPS) variants and different multi-modal encoders) and modality-centric techniques that distinguish between text-, vision-, and video-centric as well as document retrieval models. Re-ranking strategies further refine these methods via optimized example selection, relevance scoring, and filtering.
  • Fusion mechanisms are implemented through score fusion and alignment, attention-based techniques, and unified frameworks that project multimodal information into common representations.
@@
  • Generation methods include in-context learning, reasoning, instruction tuning, and source attribution.
@@ Appendix B Dataset and Benchmark
Image-Text General LAION-400M Schuhmann et al. (2021) 200M image–text pairs; used for pre-training multimodal models. Image, Text LAION-400M
Conceptual-Captions (CC) Sharma et al. (2018) 15M image–caption pairs; multilingual English–German image descriptions. Image, Text Conceptual Captions
@@
Laion-5B Schuhmann et al. (2022) 5B image–text pairs used as external memory for retrieval. Image, Text LAION-5B
COCO-CN Author and Author (2018) 20,341 images for cross-lingual tagging and captioning with Chinese sentences. Image, Text COCO-CN
@@
Video-Text BDD-X Xu et al. (2018) 77 hours of driving videos with expert textual explanations; for explainable driving behavior. Video, Text BDD-X
@@
MSR-VTT Xu et al. (2016) 10,000 videos with 20 captions each; a large video description dataset. Video, Text MSR-VTT
@@
Breakfast Kuehne et al. (2014) 1,712 videos of breakfast preparation; one of the largest fully annotated video datasets. Video, Text Breakfast
@@
MSVD-QA Xu et al. (2017) 1,970 video clips with approximately 50.5K QA pairs; video QA dataset. Video, Text MSVD-QA
@@
EpicKitchens-100 Dima (2020) 700 videos (100 hours of cooking activities) for online action prediction; egocentric vision dataset. Video, Text EPIC-KITCHENS-100
@@
Charades Sigurdsson et al. (2016) 9,848 video clips with textual descriptions; a multimodal research dataset. Video, Text Charades
@@
AudioCap Kim et al. (2019) 46K audio clips paired with human-written text captions. Audio, Text AudioCaps
@@
FashionGen Hadi Kiapour et al. (2018) 260.5K image–text pairs of fashion images and item descriptions. Image, Text Fashion-Gen
@@
Fashionpedia Author and Author (2023a) 48,000 fashion images annotated with segmentation masks and fine-grained attributes. Image, Text Fashionpedia
@@
3D ShapeNet Chang et al. (2015) 7,500 text–3D data pairs; repository for 3D CAD models. Text, 3D ShapeNet
@@
ViQuAE Biten et al. (2022) 11.8M passages from Wikipedia covering 2,397 unique entities; knowledge–intensive QA. Text ViQuAE
@@
WebQA Li et al. (2022b) 46K queries that require reasoning across text and images. Text, Image WebQA
Infoseek Li et al. (2021) Fine-grained visual knowledge retrieval using a Wikipedia–based knowledge base ( 6M passages). Image, Text Infoseek
ClueWeb22 Callan et al. (2022) 10 billion web pages organized into three subsets; a large–scale web corpus. Text ClueWeb22
@@
Other ImageNet Deng et al. (2009) 14,197,122 images for perspective understanding; a hierarchical image database. Image ImageNet
Oxford Flowers102 Nilsback and Zisserman (2008) 102 flower categories with five examples per category; image classification dataset. Image Oxford Flowers102
Stanford Cars Krause et al. (2013) Images of different car models (five examples per model); for fine-grained categorization. Image Stanford Cars
GeoDE Author and Author (2023b) 61,940 images from 40 classes across 6 world regions; emphasizes geographic diversity in object recognition. Image GeoDE
@@
M2RAG Ma et al. (2024c) Benchmarks multimodal RAG; evaluates retrieval, multi-hop reasoning, and integration. Images + Text M2RAG
Dyn-VQA Li et al. (2024b) Focuses on dynamic retrieval, multi-hop reasoning, and robustness to changing information. Images + Text Dyn-VQA
MMBench Liu et al. (2025) Covers VQA, captioning, retrieval; evaluates cross-modal understanding across vision, text, and audio. Images + Text + Audio MMBench
ScienceQA Saikh et al. (2022) Contains 21,208 questions; tests scientific reasoning with text, diagrams, and images. Images + Diagrams + Text ScienceQA
@@
Appendix C Metrics
Accuracy: Accuracy is typically defined as the ratio of correctly predicted instances to the total instances. In retrieval-based tasks, Top-K Accuracy is defined as:

### v3

@@ 4 Open Problems and Future Directions
Additional challenges and future directions about long-context processing, scalability, efficiency, and personalization are discussed in Appendix (§F).
Generalization, Explainability, and Robustness
Multimodal RAG systems often struggle with domain adaptation and exhibit modality biases, frequently over-relying on text for both retrieval and generation Winterbottom et al. (2020). Explainability remains a major challenge, as these systems often attribute responses to broad sources, citing entire documents or large visual regions instead of pinpointing exact contributing elements across modalities Ma et al. (2024b); Hu et al. (2023). Moreover, the interplay between modalities affects the outcome quality; for example, answers derived solely from text sources may differ in quality compared to those requiring a combination of text and image inputs Baltrusaitis et al. (2019). They are also vulnerable to adversarial perturbations, such as misleading images influencing textual outputs, and their performance degrades when relying on low-quality or outdated sources Chen et al. (2022b). MM-PoisonRAG Ha et al. (2025) and Poisoned-MRAG Liu et al. (2025b) demonstrate that even a few adversarial knowledge injections can hijack cross-modal retrieval and derail generation, underscoring the imperative for robust defense mechanisms against knowledge poisoning in multimodal RAG systems. While the trustworthiness of unimodal RAGs has been studied Zhou et al. (2024d), ensuring robustness in multimodal RAGs remains an open challenge and a crucial research direction.
Reasoning, Alignment, and Retrieval Enhancement
Multimodal RAGs struggle with compositional reasoning, requiring logical integration of information across modalities for coherent, context-rich outputs. While cross-modal techniques like Multimodal-CoT Zhang et al. (2023b) have emerged, further advancements are needed to enhance coherence and contextual relevance. Improving modality alignment and entity-aware retrieval is crucial. Moreover, despite the potential of knowledge graphs to enrich cross-modal reasoning, they remain underexplored in multimodal RAGs compared to text-based RAGs Zhang et al. (2024f); Procko and Ochoa (2024). Retrieval biases such as position sensitivity Hu et al. (2024c), redundancy Nan et al. (2024b), and biases from training data or retrieved content Zhai (2024), pose significant challenges. A promising direction is a unified embedding space for all modalities, enabling direct multimodal search without intermediary models (e.g., ASRs). Despite progress, mapping multimodal knowledge into a unified space remains an open challenge with substantial potential.
Agent-Based and Self-Guided Systems
Recent trends indicate a shift towards agent-based multimodal RAGs that integrate retrieval, reasoning, and generation across diverse domains. Unlike static RAGs, future systems should incorporate interactive feedback and self-guided decision-making to iteratively refine outputs. Existing feedback mechanisms often fail to determine whether errors stem from retrieval, generation, or other stages Dong et al. (2024b). The incorporation of reinforcement learning and end-to-end human-aligned feedback remains largely overlooked but holds significant potential for assessing whether retrieval is necessary, evaluating the relevance of retrieved content, and dynamically determining the most suitable modalities for response generation. Robust support for any-to-any modality is crucial for open-ended tasks Wu et al. (2024b). Future multimodal RAGs should incorporate data from diverse real-world sources, such as environmental sensors, alongside traditional modalities to enhance situational awareness. This progression aligns with the trend toward embodied AI, where models integrate knowledge with physical interaction, enabling applications in robotics, navigation, and physics-informed reasoning. Bridging retrieval-based reasoning with real-world agency brings these systems closer to AGI.
@@ 5 Conclusion
This study provides a comprehensive review of multimodal RAG, categorizing key advancements in retrieval, multimodal fusion, augmentation, generation, training strategies, and agents. We also examine task-specific applications, datasets, benchmarks, and evaluation methods while highlighting open challenges and promising future directions. We hope this work inspires future research, particularly in enhancing cross-modal reasoning and retrieval, developing agent-based interactive systems, and advancing unified multimodal embedding spaces.
@@ Appendix A Taxonomy
Figure 2 provides an overview of recent advances in multimodal RAG systems. The taxonomy is organized into several key categories.
  • Retrieval strategies cover efficient search and similarity retrieval methods (including maximum inner product search (MIPS) variants and different multimodal encoders) and modality-centric techniques that distinguish between text-, vision-, audio-, and video-centric as well as document retrieval models. Re-ranking strategies further refine these methods via optimized example selection, relevance scoring, and filtering.
  • Fusion mechanisms cover score fusion and alignment techniques, including CLIP score fusion and prototype-based embeddings that unify multimodal representations, attention-based methods such as cross-attention and co-attention transformers that dynamically weight cross-modal interactions, and unified frameworks and projections like hierarchical fusion and dense-to-sparse projections that consolidate multimodal inputs.
@@
  • Generation methods include in-context learning, reasoning, instruction tuning, source attribution, and agentic frameworks.
@@ Appendix B Dataset and Benchmark
Multimodal RAG research employs diverse datasets and benchmarks to evaluate retrieval, integration, and generation across heterogeneous sources. Image–text tasks, including captioning and retrieval, commonly use MS-COCO Lin et al. (2014), Flickr30K Young et al. (2014), and LAION-400M Schuhmann et al. (2021), while visual question answering (QA) with external knowledge is supported by OK-VQA Marino et al. (2019) and WebQA Chang et al. (2022). For complex multimodal reasoning, MultimodalQA Talmor et al. (2021) integrates text, images, and tables, whereas video-text tasks leverage ActivityNet Caba Heilbron et al. (2015) and YouCook2 Zhou et al. (2018). In the medical domain, MIMIC-CXR Johnson et al. (2019) and CheXpert Irvin et al. (2019) facilitate tasks such as medical report generation. It should be noted that a number of these datasets are unimodal (e.g., solely text-based or image-based). Unimodal datasets are frequently employed to represent a specific modality and are subsequently integrated with complementary datasets from other modalities. This modular approach allows each dataset to contribute its domain-specific strengths, thereby enhancing the overall performance of the multimodal retrieval and generation processes.
Benchmarks assess multimodal RAG systems on visual reasoning, external knowledge integration, and dynamic retrieval. The $M^{2}RAG$ Ma et al. (2024d) benchmark provides a unified evaluation framework that combines fine-grained text-modal and multimodal metrics to jointly assess both the quality of generated language and the effective integration of visual elements. In addition, Liu et al. (2025d) introduce another specialized benchmark for multimodal RAG that evaluates performance across image captioning, multi-modal question answering, fact verification, and image reranking in an open-domain retrieval setting. Vision-focused evaluations, including MRAG-Bench Hu et al. (2024c), VQAv2 Goyal et al. (2017a) and VisDoMBench Suri et al. (2024), test models on complex visual tasks. Dyn-VQA Li et al. (2024d), MMBench Liu et al. (2025c), and ScienceQA Lu et al. (2022) evaluate dynamic retrieval and multi-hop reasoning across textual, visual, and diagrammatic inputs. Knowledge-intensive benchmarks, such as TriviaQA Joshi et al. (2017) and Natural Questions Kwiatkowski et al. (2019), together with document-oriented evaluations such as OmniDocBench Ouyang et al. (2024), measure integration of unstructured and structured data. Advanced retrieval benchmarks such as RAG-Check Mortaheb et al. (2025a) evaluate retrieval relevance and system reliability, while specialized assessments like Counterfactual VQA Niu et al. (2021) test robustness against adversarial inputs. Additionally, OCR impact studies such as OHRBench Zhang et al. (2024d) examine the cascading effects of errors on RAG systems.
The choice of dataset significantly influences the evaluation focus, ranging from foundational pre-training on large-scale image-text corpora like LAION-5B Schuhmann et al. (2022) (5.85 billion pairs) or MINT-1T Awadalla et al. (2024) (3.4 billion images with 1 trillion text tokens), to more specialized tasks such as video understanding with HowTo100M Miech et al. (2019) (136 million video clips) or medical report generation using MIMIC-CXR Johnson et al. (2019) (125,417 image-report pairs).
Datasets are often tailored for specific downstream tasks. For visual question answering, VQA Antol et al. (2015) and A-OKVQA Schwenk et al. (2022) specifically require external knowledge, making them suitable for evaluating RAG systems’ ability to retrieve and reason over such knowledge. For document understanding, datasets such as DocVQA Mathew et al. (2021) and M3DocVQA Cho et al. (2024) are essential. As discussed in the benchmarks overview above, unified evaluation frameworks like $M^{2}RAG$ Ma et al. (2024d) provide a comprehensive assessment across multiple tasks, including image captioning, visual question answering, and fact verification.
Evaluating complex reasoning capabilities in multimodal RAG systems has become increasingly important. Datasets such as MultimodalQA Talmor et al. (2021), WebQA Chang et al. (2022), and ScienceQA Lu et al. (2022) are specifically designed to benchmark multi-hop reasoning abilities crucial for advanced RAG systems, with Dyn-VQA Li et al. (2024d) additionally focusing on robustness to changing information.
Comparative Analysis of Datasets
Understanding the strategic trade-offs in dataset design is crucial for multimodal RAG development, as different dataset characteristics serve distinct purposes across the model development pipeline.
(i) Scale and Diversity vs. Curation: Large-scale datasets such as LAION-5B Schuhmann et al. (2022) and Conceptual Captions Sharma et al. (2018) provide substantial scale essential for pre-training, enabling models to learn generalizable representations across diverse domains. However, their reliance on web-crawled data introduces inherent noise that can compromise training quality. Conversely, smaller, meticulously curated datasets like Flickr30K Young et al. (2014) (31,000 images with human annotations) and domain-specific collections such as Fashionpedia Jia et al. (2020) (48,000 images with segmentation masks) prioritize annotation quality over scale, making them essential for fine-tuning models and assessing specialized performance.
(ii) Modality Focus and Combination: While many systems aggregate unimodal datasets to construct multimodal contexts, datasets explicitly designed for multimodal tasks demonstrate superior alignment between modalities. Foundational datasets like MS-COCO Lin et al. (2014) and VQA Antol et al. (2015) establish benchmarks for image-text understanding, while specialized collections such as AudioSet Gemmeke et al. (2017) (2 million audio clips) and AudioCaps Kim et al. (2019) (46,000 audio clips with captions) address audio-language integration. Emerging modalities like 3D (e.g., ShapeNet Chang et al. (2015)) remain underrepresented, yet are essential for expanding RAG applications into spatial reasoning domains.
@@
Limitations of Existing Datasets and Benchmarks
While the datasets and benchmarks discussed above have significantly advanced multimodal RAG research, several limitations persist that offer important avenues for future work:
(i) Bias and Fairness: Large datasets, especially those scraped from the web, can inherit societal biases related to gender, race, or culture. This can lead to skewed model behavior and unfair outcomes. Efforts to create more balanced datasets are crucial, but comprehensive bias auditing across modalities remains a challenge.
(ii) Annotation Quality and Noise: The trade-off between dataset scale and annotation quality remains a persistent challenge. While large datasets facilitate broad learning, their often noisy or weakly supervised labels (e.g., alt-text for images) can hinder precise model training. As demonstrated by OHRBench Zhang et al. (2024d), OCR errors exemplify how noise in one modality can cascade and affect overall RAG system performance.
(iii) Coverage and Generalization Gaps: Many datasets are domain-specific, which can limit the generalization of models to out-of-domain scenarios. There is a need for more datasets covering a wider array of real-world contexts and less common modalities.
(iv) Real-World Complexity and Long-Context Understanding: Current datasets inadequately capture real-world multimodal information complexity. Challenges include efficient sampling of relevant video frames, handling multi-page documents with numerous images, and processing dynamic information environments; benchmarks like Dyn-VQA Li et al. (2024d) are, however, beginning to address this latter challenge.
(v) Lack of Adversarial and Robustness Testing: While benchmarks like Counterfactual VQA Niu et al. (2021) specifically test robustness against certain perturbations, there is a general scarcity of datasets containing multimodal adversarial examples or structured negative instances. Such datasets are vital for developing more robust and reliable RAG systems that can handle out-of-distribution inputs or misleading information.
(vi) Retrieval-Generation Integration: Many benchmarks evaluate retrieval and generation components separately rather than assessing their synergistic interplay. More holistic evaluation frameworks are needed that jointly measure retrieval accuracy, relevance of retrieved multimodal context, and final output quality, as aimed by benchmarks like MRAG-Bench Hu et al. (2024c) for visual integration and RAG-Check Mortaheb et al. (2025a) for retrieval relevance.
(vii) Limited Support for "Any-to-Any" Modalities: While current research primarily focuses on text, image, video, and audio, future RAG systems are envisioned to support any-to-any modality interactions. Existing datasets offer limited support for such comprehensive multimodality.
@@
Image-Text General LAION-400M Schuhmann et al. (2021) 400M image–text pairs; used for pre-training multimodal models. Image, Text LAION-400M
Conceptual-Captions (CC) Sharma et al. (2018) More than 3M image–caption pairs; multilingual English–German image descriptions. Image, Text Conceptual Captions
@@
Laion-5B Schuhmann et al. (2022) 5.85B image–text pairs used as external memory for retrieval. Image, Text LAION-5B
COCO-CN Li et al. (2019) 20,341 images for cross-lingual tagging and captioning with Chinese sentences. Image, Text COCO-CN
@@
MINT-1T Awadalla et al. (2024) 1T text tokens and 3.4B images; 10x larger than existing open-source datasets. Image, Text MINT-1T
ShareGPT4V Chen et al. (2024c) 1.2M images with GPT-4-generated captions, including spatial and factual details. Image, Text ShareGPT4V
OmniCorpus Li et al. (2025b) 8.6B images and 1.7T tokens across 2.2B web documents; interleaved text-image layout. Image, Text OmniCorpus
Video-Text BDD-X Kim et al. (2018) 77 hours of driving videos with expert textual explanations; for explainable driving behavior. Video, Text BDD-X
@@
EpicKitchens-100 Damen et al. (2022) 700 videos (100 hours of cooking activities) for online action prediction; egocentric vision dataset. Video, Text EPIC-KITCHENS-100
@@
InternVid Wang et al. (2023b) 7M YouTube videos (760K hours), 234M clips, 4.1B words; used for video-text pretraining and representation learning. Video, Text InternVid
OpenVid-1M Nan et al. (2024a) 1 million video-text pairs for multimodal learning. Video, Text OpenVid-1M
@@
AudioCaps Kim et al. (2019) 46K audio clips paired with human-written text captions. Audio, Text AudioCaps
MusicCaps Agostinelli et al. (2023) It is composed of 5.5k music-text pairs, with rich text descriptions provided by human experts. Audio, Text MusicCaps
Clotho Drossos et al. (2020) Audio captioning dataset with diverse soundscapes. Audio, Text Clotho
WavCaps Mei et al. (2024) Large-scale weakly-labeled audio-text dataset, comprising approximately 400k audio clips with paired captions. Audio, Text WavCaps
Spoken SQuAD Li et al. (2018) Audio version of the SQuAD dataset for spoken question answering, focusing on the listening comprehension task. Audio, Text Spoken SQuAD
@@
Quilt-1M Ikezogwo et al. (2023) 438K medical images with 768K text pairs; includes microscopic images and UMLS entities. Image, Text Quilt-1M
@@
FashionGen Rostamzadeh et al. (2018) 260.5K image–text pairs of fashion images and item descriptions. Image, Text Fashion-Gen
@@
Fashionpedia Jia et al. (2020) 48,000 fashion images annotated with segmentation masks and fine-grained attributes. Image, Text Fashionpedia
@@
3D ShapeNet Chang et al. (2015) Covering 55 common object categories with  51,300 unique 3D models. Text, 3D ShapeNet
@@
MultimodalQA Talmor et al. (2021) 29,918 questions requiring multi-modal multi-hop reasoning over text, tables, and images. Image, Text, Table MultimodalQA
ViQuAE Lerner et al. (2022) 11.8M passages from Wikipedia covering 2,397 unique entities; knowledge–intensive QA. Text ViQuAE
@@
WebQA Chang et al. (2022) 46K queries that require reasoning across text and images. Text, Image WebQA
Infoseek Chen et al. (2023) Fine-grained visual knowledge retrieval using a Wikipedia–based knowledge base ( 6M passages). Image, Text Infoseek
ClueWeb22 Overwijk et al. (2022) 10 billion web pages organized into three subsets; a large–scale web corpus. Text ClueWeb22
@@
DocVQA Mathew et al. (2021) 12,767 diverse document images with 50K QA pairs, categorized by reasoning type to evaluate DocVQA methods. Image, Text DocVQA
ChartQA Masry et al. (2022) 9.6K human-written QA pairs + 23.1K generated from chart summaries. Image, Text ChartQA
DVQA Kafle et al. (2018) 3.5M QA pairs on 300K diagrams, evaluating structure, data retrieval, and reasoning. Image, Text DVQA
RETVQA Penamakuri et al. (2023) 416,000 QA samples where retrieval from a large image set is needed to answer questions; emphasizes RAG pipeline. Image, Text RETVQA
@@
M3DocVQA Cho et al. (2024) 2,441 multi-hop questions across 3,368 PDF documents; evaluates open-domain DocVQA. Image, Text M3DocVQA
MMLongBench-Doc Ma et al. (2024c) 135 lengthy PDFs with 1,091 questions; focuses on multi-hop reasoning in single documents. Image, Text MMLongBench-Doc
Other GeoDE Ramaswamy et al. (2023) 61,940 images from 40 classes across 6 world regions; emphasizes geographic diversity in object recognition. Image GeoDE
RU-AI Huang et al. (2025) 1.47M samples of real vs AI-generated content for fake detection robustness. Image, Text, Audio RU-AI
MIMIC-IT Li et al. (2025a) 2.8M multimodal instruction-response pairs for model alignment. Image, Video, Text MIMIC-IT
MMVQA Ding et al. (2024c) 262K question-answer pairs across 3,146 multipage research PDFs for robust multimodal information retrieval. Image, Text MMVQA
@@
$M^{2}RAG$ Ma et al. (2024d) Benchmarks multimodal RAG; evaluates retrieval, multi-hop reasoning, and integration. Images + Text $M^{2}RAG$
Dyn-VQA Li et al. (2024d) Focuses on dynamic retrieval, multi-hop reasoning, and robustness to changing information. Images + Text Dyn-VQA
MMBench Liu et al. (2025c) Covers VQA, captioning, retrieval; evaluates cross-modal understanding across vision, text, and audio. Images + Text + Audio MMBench
ScienceQA Lu et al. (2022) Contains 21,208 questions; tests scientific reasoning with text, diagrams, and images. Images + Diagrams + Text ScienceQA
@@
{forest}
forked edges, for tree= child anchor=west, parent anchor=east, grow’=east, anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, minimum height=2em, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=0.4em, inner ysep=0.6em, line width=0.8pt, text width=8.5em, ver/.style= rotate=90, child anchor=north, parent anchor=south, anchor=center, text width=17em, font=, minimum height=3em, minimum width=5em, , leaf/.style= fill=green!20, text= black, text width=44.5em, font=, inner xsep=0.4em, inner ysep=0.6em, draw, , intermediate/.style= font=, minimum height=2.5em, minimum width=18em, text width=18em, draw, , [ Multimodal RAG Application Domains, ver , fill=white!20 [ Healthcare and Medicine (§E) , fill=yellow!20, intermediate [ MMED-RAG Xia et al. (2024a), RULE Xia et al. (2024b), AsthmaBot Bahaj and Ghogho (2024), Realm Zhu et al. (2024c),  Su et al. (2024a), FactMM-RAG Sun et al. (2024b), RA-RRG Choi et al. (2025) , leaf ] ] [ Software Engineering (§E) , fill=yellow!20, intermediate [ DocPrompting Zhou et al. (2023), RACE Shi et al. (2022), CEDAR Nashid et al. (2023), REDCODER Parvez et al. (2021) , leaf ] ] [ Fashion and E-Commerce (§E) , fill=yellow!20, intermediate [ Unifashion Zhao et al. (2024),  Dang (2024), Fashion-RAG Sanguigni et al. (2025), LLM4DESIGN Chen et al. (2024d) , leaf ] ] [ Entertainment and Social Computing (§E) , fill=yellow!20, intermediate [ SoccerRAG Strand et al. (2024), MMRA Zhong et al. (2024) , leaf ] ] [ Emerging Applications (§E) , fill=yellow!20, intermediate [ RAG-Driver Yuan et al. (2024), ENWAR Nazar et al. (2024),  Riedler and Langer (2024), Img2Loc Zhou et al. (2024e) , leaf ] ] ]
Figure 3: Taxonomy of application domains for Multimodal Retrieval-Augmented Generation systems.
Appendix C Evaluation and Metrics
Evaluating multimodal RAG models is complex due to their varied input types and complex structure. The evaluation combines metrics from VLMs, generative AI, and retrieval systems to assess capabilities like text/image generation and information retrieval. Our review found about 60 different metrics used in the field. In the following paragraphs, we will examine the most important and widely used metrics for evaluating multimodal RAG.
Retrieval Evaluation
Retrieval performance is measured through accuracy, recall, and precision metrics, with an F1 score combining recall and precision. Accuracy is typically defined as the ratio of correctly predicted instances to the total instances. In retrieval-based tasks, Top-K Accuracy is defined as: